OpenStudy (anonymous):

Simplify 6x+24 over 6. 6x + 4 x + 24 x x + 4

OpenStudy (whpalmer4):

Divide each term by 6, what do you get?

OpenStudy (anonymous):

x+4? cuz the 6 goes away right? and the 24 divided by 6 ?!

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Exactly! What you are doing is the reverse of 6(x+4) = 6x + 24 and then you end up with \[\frac{6(x+4)}{6} = (x+4)\] after the 6s cancel out.

OpenStudy (anonymous):

omg thanks!! Im getting better at this :)))